# Kevin Wang

### v3

@@ 1 Introduction
While scaling model size has been an effective recipe in many areas of machine learning, its role and impact in reinforcement learning (RL) remains unclear. The typical model size for state-based RL tasks is between 2 to 5 layers (Raffin et al., , 2021; Huang et al., , 2022). In contrast, it is not uncommon to use very deep networks in other domain areas; Llama 3 (Dubey et al., , 2024) and Stable Diffusion 3 (Esser et al., , 2024) have hundreds of layers. In fields such as vision (Radford et al., , 2021; Zhai et al., , 2021; Dehghani et al., , 2023) and language (Srivastava et al., , 2023), models often only acquire the ability to solve certain tasks once the model is larger than a critical scale. In the RL setting, many researchers have searched for similar emergent phenomena (Srivastava et al., , 2023), but these papers typically report only small marginal benefits and typically only on tasks where small models already achieve some degree of success (Nauman et al., 2024b, ; Lee et al., , 2024; Farebrother et al., , 2024). A key open question in RL today is whether it is possible to achieve similar jumps in performance by scaling RL networks.
@@
At first glance, it makes sense why training very large RL networks should be difficult: the RL problem provides very few bits of feedback (e.g., only a sparse reward after a long sequence of observations), so the ratio of feedback to parameters is very small. The conventional wisdom (LeCun, , 2016), reflected in many recent models (Radford, , 2018; Chen et al., , 2020; Goyal et al., , 2019), has been that large AI systems must be trained primarily in a self-supervised fashion and that RL should only be used to finetune these models. Indeed, many of the recent breakthroughs in other fields have been primarily achieved with self-supervised methods, whether in computer vision (Caron et al., , 2021; Radford et al., , 2021; Liu et al., , 2024), NLP (Srivastava et al., , 2023) or multimodal learning (Zong et al., , 2024). Thus, if we hope to scale reinforcement learning methods, self-supervision will likely be a key ingredient.
@@
In this paper, we will study building blocks for scaling reinforcement learning. Our first step is to rethink the conventional wisdom above: “reinforcement learning” and “self-supervised learning” are not diametric learning rules, but rather can be married together into self-supervised RL systems that explore and learn policies without reference to a reward function or demonstrations (Eysenbach et al., , 2021, 2022; Lee et al., , 2022). In this work, we use one of the simplest self-supervised RL algorithms, contrastive RL (CRL) (Eysenbach et al., , 2022). The second step is to recognize the importance of increasing available data. We will do this by building on recent GPU-accelerated RL frameworks (Makoviychuk et al., , 2021; Rutherford et al., , 2023; Rudin et al., , 2022; Bortkiewicz et al., , 2024). The third step is to increase network depth, using networks that are up to $100\times$ deeper than those typical in prior work. Stabilizing the training of such networks will require incorporating architectural techniques from prior work, including residual connections (He et al., , 2015), layer normalization (Ba et al., , 2016), and Swish activation (Ramachandran et al., , 2018). Our experiments will also study the relative importance of batch size and network width.
@@
  • Empirical Scalability: We observe a significant performance increase, more than $20\times$ in half of the environments and out-performing other standard goal-conditioned baselines. These performances gains correspond to qualitatively distinct policies that emerge with scale.
  • Scaling Depth in Network Architecture: While many prior RL works have primarily focused on increasing network width, they often report limited or even negative returns when expanding depth (Lee et al., , 2024; Nauman et al., 2024b, ). In contrast, our approach unlocks the ability to scale along the axis of depth, yielding performance improvements that surpass those from scaling width alone (see Sec. 4).
@@ 2 Related Work
Natural Language Processing (NLP) and Computer Vision (CV) have recently converged in adopting similar architectures (i.e. transformers) and shared learning paradigms (i.e self-supervised learning), which together have enabled transformative capabilities of large-scale models (Vaswani et al., , 2017; Srivastava et al., , 2023; Zhai et al., , 2021; Dehghani et al., , 2023; Wei et al., , 2022). In contrast, achieving similar advancements in reinforcement learning (RL) remains challenging. Several studies have explored the obstacles to scaling large RL models, including parameter underutilization (Obando-Ceron et al., , 2024), plasticity and capacity loss (Lyle et al., , 2024, 2022), data sparsity (Andrychowicz et al., , 2017; LeCun, , 2016), and training instabilities (Ota et al., , 2021; Henderson et al., , 2018; Van Hasselt et al., , 2018; Nauman et al., 2024a, ). As a result, current efforts to scale RL models are largely restricted to specific problem domains, such as imitation learning (Tuyls et al., , 2024), multi-agent games (Neumann and Gros, , 2022), language-guided RL (Driess et al., , 2023; Ahn et al., , 2022), and discrete action spaces (Obando-Ceron et al., , 2024; Schwarzer et al., , 2023).
@@
Recent approaches suggest several promising directions, including new architectural paradigms (Obando-Ceron et al., , 2024), distributed training approaches (Ota et al., , 2021; Espeholt et al., , 2018), distributional RL (Kumar et al., , 2023), and distillation (Team et al., , 2023). Compared to these approaches, our method makes a simple extension to an existing self-supervised RL algorithm. The most recent works in this vein include Lee et al., (2024) and Nauman et al., 2024b , which leverage residual connections to facilitate the training of wider networks. These efforts primarily focus on network width, noting limited gains from additional depth, thus both works use architectures with only four MLP layers. In our method, we find that scaling width indeed improves performance (Section˜4.4); however, our approach also enables scaling along depth, proving to be more powerful than width alone.
@@
One notable effort to train deeper networks is described by Farebrother et al., (2024), who cast value-based RL into a classification problem by discretizing the TD objective into a categorical cross-entropy loss. This approach draws on the conjecture that classification-based methods can be more robust and stable and thus may exhibit better scaling properties than their regressive counterparts (Torgo and Gama, , 1996; Farebrother et al., , 2024). The CRL algorithm that we use effectively uses a cross-entropy loss as well (Eysenbach et al., , 2022). Its InfoNCE objective is a generalization of the cross-entropy loss, thereby performing RL tasks by effectively classifying whether current states and actions belong to the same or different trajectory that leads toward a goal state. In this vein, our work serves as a second piece of evidence that classification, much like cross-entropy’s role in the scaling success in NLP, could be a potential building block in RL.
@@ 3 Preliminaries
We define a goal-conditioned MDP as tuple $\mathcal{M}_{g}=(\mathcal{S},\mathcal{A},p_{0},p,p_{g},r_{g},\gamma)$, where the agent interacts with the environment to reach arbitrary goals (Kaelbling, , 1993; Andrychowicz et al., , 2017; Blier et al., , 2021). At every time step $t$, the agent observes state $s_{t}\in\mathcal{S}$ and performs a corresponding action $a_{t}\in\mathcal{A}$. The agent starts interaction in states sampled from $p_{0}(s_{0})$, and the interaction dynamics are defined by the transition probability distribution $p(s_{t+1}\mid s_{t},a_{t})$. Goals $g\in\mathcal{G}$ are defined in a goal space $\mathcal{G}$, which is related to $\mathcal{S}$ via a mapping $f:\mathcal{S}\to\mathcal{G}$. For example, $\mathcal{G}$ may correspond to a subset of state dimensions. The prior distribution over goals is defined by $p_{g}(g)$. The reward function is defined as the probability density of reaching the goal in the next time step $r_{g}(s_{t},a_{t})\triangleq(1-\gamma)p(s_{t+1}=g\mid s_{t},a_{t})$, with discount factor $\gamma$.
@@
Our experiments will use the contrastive RL algorithm (Eysenbach et al., , 2022) to solve goal-conditioned problems. Contrastive RL is an actor-critic method; we will use $f_{\phi,\psi}(s,a,g)$ to denote the critic and $\pi_{\theta}(a\mid s,g)$ to denote the policy. The critic is parametrized with two neural networks that return state, action pair embedding $\phi(s,a)$ and goal embedding $\psi(g)$. The critic’s output is defined as the $l^{2}$-norm between these embeddings: $f_{\phi,\psi}(s,a,g)=\|\phi(s,a)-\psi(g)\|_{2}$. The critic is trained with the InfoNCE objective (Sohn, , 2016) as in previous works (Eysenbach et al., , 2022, 2021; Zheng et al., , 2023, 2024; Myers et al., , 2024; Bortkiewicz et al., , 2024). Training is conducted on batches $\mathcal{B}$, where $\color[rgb]{0,0.55,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.55,0}s_{i},a_{i},g_{i}$} represent the state, action, and goal (future state) sampled from the same trajectory, while $\color[rgb]{0.75,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.75,0,0}g_{j}$} represents a goal sampled from a different, random trajectory. The objective function is defined as:
@@
We incorporate residual connections (He et al., , 2015) into our architecture, following their successful use in RL (Farebrother et al., , 2024; Lee et al., , 2024; Nauman et al., 2024b, ). A residual block transforms a given representation $\mathbf{h}_{i}$ by adding a learned residual function $F_{i}(\mathbf{h}_{i})$ to the original representation. Mathematically, this is expressed as:
@@
where $\mathbf{h}_{i+1}$ is the output representation, $\mathbf{h}_{i}$ is the input representation, and $F_{i}(\mathbf{h}_{i})$ is a transformation learned through the network (e.g., using one or more layers). The addition ensures that the network learns modifications to the input rather than entirely new transformations, helping to preserve useful features from earlier layers. Residual connections improve gradient propagation by introducing shortcut paths (He et al., , 2016; Veit et al., , 2016), enabling more effective training of deep models.
@@ 4.1 Experimental Setup
All RL experiments use the JaxGCRL codebase (Bortkiewicz et al., , 2024), which facilitates fast online GCRL experiments based on Brax (Freeman et al., , 2021) and MJX (Todorov et al., , 2012) environments. The specific environments used are a range of locomotion, navigation, and robotic manipulation tasks, for details see Appendix˜B. We use a sparse reward setting, with $r=1$ only when the agent is in the goal proximity. For evaluation, we measure the number of time steps (out of 1000) that the agent is near the goal. When reporting an algorithm’s performance as a single number, we compute the average score over the last five epochs of training.
@@
We employ residual connections from the ResNet architecture (He et al., , 2015), with each residual block consisting of four repeated units of a Dense layer, a Layer Normalization (Ba et al., , 2016) layer, and Swish activation (Ramachandran et al., , 2018). We apply the residual connections immediately following the final activation of the residual block, as shown in Figure˜2. In this paper, we define the depth of the network as the total number of Dense layers across all residual blocks in the architecture. In all experiments, the depth refers to the configuration of the actor network and both critic encoder networks, which are scaled jointly, except for the ablation experiment in Section˜4.4.
@@ 4.2 Scaling Depth in Contrastive RL
We start by studying how increasing network depth can increase performance. Both the JaxGCRL benchmark and relevant prior work (Lee et al., , 2024; Nauman et al., 2024b, ; Zheng et al., , 2024) use MLPs with a depth of 4, and as such we adopt it as our baseline. In contrast, we will study networks of depth 8, 16, 32, and 64. The results in Figure˜1 demonstrate that deeper networks achieve significant performance improvements across a diverse range of locomotion, navigation, and manipulation tasks. Compared to the 4-layer models typical in prior work, deeper networks achieve $2-5\times$ gains in robotic manipulation tasks, over $20\times$ gains in long-horizon maze tasks such as Ant U4-Maze and Ant U5-Maze, and over $50\times$ gains in humanoid-based tasks. The full table of performance increases up to depth 64 is provided in Table˜1.
@@ 4.4 What Matters for CRL Scaling
Figure 4: Scaling network width vs. depth. Here, we reflect findings from previous works (Lee et al., , 2024; Nauman et al., 2024b, ) which suggest that increasing network width can enhance performance. However, in contrast to prior work, our method is able to scale depth, yielding more impactful performance gains. For instance, in the Humanoid environment, raising the width to 2048 (depth=4) fails to match the performance achieved by simply doubling the depth to 8 (width=256). The comparative advantage of scaling depth is more pronounced as the observational dimensionality increases.
Past literature has shown that scaling network width can be effective  (Lee et al., , 2024; Nauman et al., 2024b, ). In Figure˜4, we find that scaling width is also helpful in our experiments: wider networks consistently outperform narrower networks (depth held constant at 4). However, depth seems to be a more effective axis for scaling: simply doubling the depth to 8 (width held constant at 256) outperforms the widest networks in all three environments. The advantage of depth scaling is most pronounced in the Humanoid environment (observation dimension 268), followed by Ant Big Maze (dimension 29) and Arm Push Easy (dimension 17), suggesting that the comparative benefit may increase with higher observation dimensionality.
@@
To investigate the role of scaling in the actor and critic networks, Figure˜6 presents the final performance for various combinations of actor and critic depths across three environments. Prior work (Nauman et al., 2024b, ; Lee et al., , 2024) focuses on scaling the critic network, finding that scaling the actor degrades performance. In contrast, while we do find that scaling the critic is more impactful in two of the three environments (Humanoid, Arm Push Easy), our method benefits from scaling the actor network jointly, with one environment (Ant Big Maze) demonstrating actor scaling to be more impactful. Thus, our method suggests that scaling both the actor and critic networks can play a complementary role in enhancing performance.
@@
Scaling batch size has been well-established in other areas of machine learning (Chen et al., , 2022; Zhang et al., , 2024). However, this approach has not translated as effectively to reinforcement learning (RL), and prior work has even reported negative impacts on value-based RL (Obando-Ceron et al., , 2023). Indeed, in our experiments, simply increasing the batch size for the original CRL networks yields only marginal differences in performance (Figure˜7, top left).
At first glance, this might seem counterintuitive: since reinforcement learning typically involves fewer informational bits per piece of training data (LeCun, , 2016), one might expect higher variance in batch loss or gradients, suggesting the need for larger batch sizes to compensate. At the same time, this possibility hinges on whether the model in question can actually make use of a bigger batch size—in domains of ML where scaling has been successful, larger batch sizes usually bring the most benefit when coupled with sufficiently large models (Zhang et al., , 2024; Chen et al., , 2022). One hypothesis is that the small models traditionally used in RL may obscure the underlying benefits of larger batch size.

### v4

@@ 1 Introduction
While scaling model size has been an effective recipe in many areas of machine learning, its role and impact in reinforcement learning (RL) remain unclear. The typical model size for state-based RL tasks is between 2 to 5 layers (Raffin et al., 2021; Huang et al., 2022). In contrast, it is not uncommon to use very deep networks in other domain areas; Llama 3 (Dubey et al., 2024) and Stable Diffusion 3 (Esser et al., 2024) have hundreds of layers. In fields such as vision (Radford et al., 2021; Zhai et al., 2021; Dehghani et al., 2023) and language (Srivastava et al., 2023), models often only acquire the ability to solve certain tasks once they are larger than a critical scale. In the RL setting, many researchers have searched for similar emergent phenomena (Srivastava et al., 2023), but these papers typically report only small marginal benefits and typically only on tasks where small models already achieve some degree of success (Nauman et al., 2024b; Lee et al., 2024; Farebrother et al., 2024). A key open question in RL today is whether it is possible to achieve similar jumps in performance by scaling RL networks.
@@
At first glance, it makes sense why training very large RL networks should be difficult: the RL problem provides very few bits of feedback (e.g., only a sparse reward after a long sequence of observations), so the ratio of feedback to parameters is very small. The conventional wisdom (LeCun, 2016), reflected in many recent models (Radford, 2018; Chen et al., 2020; Goyal et al., 2019), has been that large AI systems must be trained primarily in a self-supervised fashion and that RL should only be used to finetune these models. Indeed, many of the recent breakthroughs in other fields have been primarily achieved with self-supervised methods, whether in computer vision (Caron et al., 2021; Radford et al., 2021; Liu et al., 2024), NLP (Srivastava et al., 2023), or multimodal learning (Zong et al., 2024). Thus, if we hope to scale reinforcement learning methods, self-supervision will likely be a key ingredient.
@@
In this paper, we will study building blocks for scaling reinforcement learning. Our first step is to rethink the conventional wisdom above: “reinforcement learning” and “self-supervised learning” are not diametric learning rules, but rather can be married together into self-supervised RL systems that explore and learn policies without reference to a reward function or demonstrations (Eysenbach et al., 2021, 2022; Lee et al., 2022). In this work, we use one of the simplest self-supervised RL algorithms, contrastive RL (CRL) (Eysenbach et al., 2022). The second step is to recognize the importance of increasing data availability. We will do this by building on recent GPU-accelerated RL frameworks (Makoviychuk et al., 2021; Rutherford et al., 2023; Rudin et al., 2022; Bortkiewicz et al., 2024). The third step is to increase network depth, using networks that are up to $100\times$ deeper than those typically found in prior work. Stabilizing the training of such networks will require incorporating architectural techniques from prior work, including residual connections (He et al., 2015), layer normalization (Ba et al., 2016), and Swish activation (Ramachandran et al., 2018). Our experiments will also study the relative importance of batch size and network width.
@@
  • Empirical Scalability: We observe a significant performance increase, more than $20\times$ in half of the environments and outperforming other standard goal-conditioned baselines. These performance gains correspond to qualitatively distinct policies that emerge as the scale increases.
  • Scaling Depth in Network Architecture: While many prior RL works have primarily focused on increasing network width, they often report limited or even negative returns when expanding depth (Lee et al., 2024; Nauman et al., 2024b). In contrast, our approach unlocks the ability to scale along the axis of depth, yielding performance improvements that surpass those from scaling width alone (see Sec. 4).
@@ 2 Related Work
Natural Language Processing (NLP) and Computer Vision (CV) have recently converged in adopting similar architectures (i.e. transformers) and shared learning paradigms (i.e self-supervised learning), which together have enabled transformative capabilities of large-scale models (Vaswani et al., 2017; Srivastava et al., 2023; Zhai et al., 2021; Dehghani et al., 2023; Wei et al., 2022). In contrast, achieving similar advancements in reinforcement learning (RL) remains challenging. Several studies have explored the obstacles to scaling large RL models, including parameter underutilization (Obando-Ceron et al., 2024), plasticity and capacity loss (Lyle et al., 2024, 2022), data sparsity (Andrychowicz et al., 2017; LeCun, 2016), and training instabilities (Ota et al., 2021; Henderson et al., 2018; Van Hasselt et al., 2018; Nauman et al., 2024a). As a result, current efforts to scale RL models are largely restricted to specific problem domains, such as imitation learning (Tuyls et al., 2024), multi-agent games (Neumann and Gros, 2022), language-guided RL (Driess et al., 2023; Ahn et al., 2022), and discrete action spaces (Obando-Ceron et al., 2024; Schwarzer et al., 2023).
@@
Recent approaches suggest several promising directions, including new architectural paradigms (Obando-Ceron et al., 2024), distributed training approaches (Ota et al., 2021; Espeholt et al., 2018), distributional RL (Kumar et al., 2023), and distillation (Team et al., 2023). Compared to these approaches, our method makes a simple extension to an existing self-supervised RL algorithm. The most recent works in this vein include Lee et al. (2024) and Nauman et al. (2024b), which leverage residual connections to facilitate the training of wider networks. These efforts primarily focus on network width, noting limited gains from additional depth, thus both works use architectures with only four MLP layers. In our method, we find that scaling width indeed improves performance (Section˜4.4); however, our approach also enables scaling along depth, proving to be more powerful than width alone.
@@
One notable effort to train deeper networks is described by Farebrother et al. (2024), who cast value-based RL into a classification problem by discretizing the TD objective into a categorical cross-entropy loss. This approach draws on the conjecture that classification-based methods can be more robust and stable and thus may exhibit better scaling properties than their regressive counterparts (Torgo and Gama, 1996; Farebrother et al., 2024). The CRL algorithm that we use effectively uses a cross-entropy loss as well (Eysenbach et al., 2022). Its InfoNCE objective is a generalization of the cross-entropy loss, thereby performing RL tasks by effectively classifying whether current states and actions belong to the same or different trajectory that leads toward a goal state. In this vein, our work serves as a second piece of evidence that classification, much like cross-entropy’s role in the scaling success in NLP, could be a potential building block in RL.
@@ 3 Preliminaries
We define a goal-conditioned MDP as tuple $\mathcal{M}_{g}=(\mathcal{S},\mathcal{A},p_{0},p,p_{g},r_{g},\gamma)$, where the agent interacts with the environment to reach arbitrary goals (Kaelbling, 1993; Andrychowicz et al., 2017; Blier et al., 2021). At every time step $t$, the agent observes state $s_{t}\in\mathcal{S}$ and performs a corresponding action $a_{t}\in\mathcal{A}$. The agent starts interaction in states sampled from $p_{0}(s_{0})$, and the interaction dynamics are defined by the transition probability distribution $p(s_{t+1}\mid s_{t},a_{t})$. Goals $g\in\mathcal{G}$ are defined in a goal space $\mathcal{G}$, which is related to $\mathcal{S}$ via a mapping $f:\mathcal{S}\to\mathcal{G}$. For example, $\mathcal{G}$ may correspond to a subset of state dimensions. The prior distribution over goals is defined by $p_{g}(g)$. The reward function is defined as the probability density of reaching the goal in the next time step $r_{g}(s_{t},a_{t})\triangleq(1-\gamma)p(s_{t+1}=g\mid s_{t},a_{t})$, with discount factor $\gamma$.
@@
Our experiments will use the contrastive RL algorithm (Eysenbach et al., 2022) to solve goal-conditioned problems. Contrastive RL is an actor-critic method; we will use $f_{\phi,\psi}(s,a,g)$ to denote the critic and $\pi_{\theta}(a\mid s,g)$ to denote the policy. The critic is parametrized with two neural networks that return state, action pair embedding $\phi(s,a)$ and goal embedding $\psi(g)$. The critic’s output is defined as the $l^{2}$-norm between these embeddings: $f_{\phi,\psi}(s,a,g)=\|\phi(s,a)-\psi(g)\|_{2}$. The critic is trained with the InfoNCE objective (Sohn, 2016) as in previous works (Eysenbach et al., 2022, 2021; Zheng et al., 2023, 2024; Myers et al., 2024; Bortkiewicz et al., 2024). Training is conducted on batches $\mathcal{B}$, where $\color[rgb]{0,0.55,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.55,0}s_{i},a_{i},g_{i}$} represent the state, action, and goal (future state) sampled from the same trajectory, while $\color[rgb]{0.75,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.75,0,0}g_{j}$} represents a goal sampled from a different, random trajectory. The objective function is defined as:
@@
We incorporate residual connections (He et al., 2015) into our architecture, following their successful use in RL (Farebrother et al., 2024; Lee et al., 2024; Nauman et al., 2024b). A residual block transforms a given representation $\mathbf{h}_{i}$ by adding a learned residual function $F_{i}(\mathbf{h}_{i})$ to the original representation. Mathematically, this is expressed as:
@@
where $\mathbf{h}_{i+1}$ is the output representation, $\mathbf{h}_{i}$ is the input representation, and $F_{i}(\mathbf{h}_{i})$ is a transformation learned through the network (e.g., using one or more layers). The addition ensures that the network learns modifications to the input rather than entirely new transformations, helping to preserve useful features from earlier layers. Residual connections improve gradient propagation by introducing shortcut paths (He et al., 2016; Veit et al., 2016), enabling more effective training of deep models.
@@ 4.1 Experimental Setup
All RL experiments use the JaxGCRL codebase (Bortkiewicz et al., 2024), which facilitates fast online GCRL experiments based on Brax (Freeman et al., 2021) and MJX (Todorov et al., 2012) environments. The specific environments used are a range of locomotion, navigation, and robotic manipulation tasks, for details see Appendix˜B. We use a sparse reward setting, with $r=1$ only when the agent is in the goal proximity. For evaluation, we measure the number of time steps (out of 1000) that the agent is near the goal. When reporting an algorithm’s performance as a single number, we compute the average score over the last five epochs of training.
@@
We employ residual connections from the ResNet architecture (He et al., 2015), with each residual block consisting of four repeated units of a Dense layer, a Layer Normalization (Ba et al., 2016) layer, and Swish activation (Ramachandran et al., 2018). We apply the residual connections immediately following the final activation of the residual block, as shown in Figure˜2. In this paper, we define the depth of the network as the total number of Dense layers across all residual blocks in the architecture. In all experiments, the depth refers to the configuration of the actor network and both critic encoder networks, which are scaled jointly, except for the ablation experiment in Section˜4.4.
@@ 4.2 Scaling Depth in Contrastive RL
We start by studying how increasing network depth can increase performance. Both the JaxGCRL benchmark and relevant prior work (Lee et al., 2024; Nauman et al., 2024b; Zheng et al., 2024) use MLPs with a depth of 4, and as such we adopt it as our baseline. In contrast, we will study networks of depth 8, 16, 32, and 64. The results in Figure˜1 demonstrate that deeper networks achieve significant performance improvements across a diverse range of locomotion, navigation, and manipulation tasks. Compared to the 4-layer models typical in prior work, deeper networks achieve $2-5\times$ gains in robotic manipulation tasks, over $20\times$ gains in long-horizon maze tasks such as Ant U4-Maze and Ant U5-Maze, and over $50\times$ gains in humanoid-based tasks. The full table of performance increases up to depth 64 is provided in Table˜1.
@@ 4.4 What Matters for CRL Scaling
Figure 4: Scaling network width vs. depth. Here, we reflect findings from previous works (Lee et al., 2024; Nauman et al., 2024b) which suggest that increasing network width can enhance performance. However, in contrast to prior work, our method is able to scale depth, yielding more impactful performance gains. For instance, in the Humanoid environment, raising the width to 2048 (depth=4) fails to match the performance achieved by simply doubling the depth to 8 (width=256). The comparative advantage of scaling depth is more pronounced as the observational dimensionality increases.
Past literature has shown that scaling network width can be effective  (Lee et al., 2024; Nauman et al., 2024b). In Figure˜4, we find that scaling width is also helpful in our experiments: wider networks consistently outperform narrower networks (depth held constant at 4). However, depth seems to be a more effective axis for scaling: simply doubling the depth to 8 (width held constant at 256) outperforms the widest networks in all three environments. The advantage of depth scaling is most pronounced in the Humanoid environment (observation dimension 268), followed by Ant Big Maze (dimension 29) and Arm Push Easy (dimension 17), suggesting that the comparative benefit may increase with higher observation dimensionality.
@@
To investigate the role of scaling in the actor and critic networks, Figure˜6 presents the final performance for various combinations of actor and critic depths across three environments. Prior work (Nauman et al., 2024b; Lee et al., 2024) focuses on scaling the critic network, finding that scaling the actor degrades performance. In contrast, while we do find that scaling the critic is more impactful in two of the three environments (Humanoid, Arm Push Easy), our method benefits from scaling the actor network jointly, with one environment (Ant Big Maze) demonstrating actor scaling to be more impactful. Thus, our method suggests that scaling both the actor and critic networks can play a complementary role in enhancing performance.
@@
Scaling batch size has been well-established in other areas of machine learning (Chen et al., 2022; Zhang et al., 2024). However, this approach has not translated as effectively to reinforcement learning (RL), and prior work has even reported negative impacts on value-based RL (Obando-Ceron et al., 2023). Indeed, in our experiments, simply increasing the batch size for the original CRL networks yields only marginal differences in performance (Figure˜7, top left).
At first glance, this might seem counterintuitive: since reinforcement learning typically involves fewer informational bits per piece of training data (LeCun, 2016), one might expect higher variance in batch loss or gradients, suggesting the need for larger batch sizes to compensate. At the same time, this possibility hinges on whether the model in question can actually make use of a bigger batch size—in domains of ML where scaling has been successful, larger batch sizes usually bring the most benefit when coupled with sufficiently large models (Zhang et al., 2024; Chen et al., 2022). One hypothesis is that the small models traditionally used in RL may obscure the underlying benefits of larger batch size.